how many photons are produced in a laser pulse od 0.301 J at 655 nm?
Answer #1

Answer

9.92 × 1017 photons

Explanation

Energy of photon , E = h c /\lambda

where,

h = planck constant , 6.626 ×10-34 J.s

c = speed of light , 2.998×108m/s

  \lambda = wavelength of photon , 655nm = 655 ×10-9m

E = 6.626 ×10-34 J.s ( 2.998 × 108m/s/655×10-9m)

E = 3.033 × 10-19J

Number of photons produced = 0.301J/3.033×10-19J /photon = 9.92 × 1017 photons